# Review on the decoding algorithms for surface codes

### v4

@@ 5.3.1 BP: Specifics
Earlier we introduced QMLD and DQMLD as the two possible domains of QECC decoding problems. We also mentioned how QMLD and DQMLD are both intractable problems (QMLD belongs to the NP-complete complexity class while DQMLD belongs to the #P-complete complexity class). This means that decoding algorithms generally work by finding solutions to good aproximations of the decoding problem. This is precisely the operating principle behind BP works. In the classical context, BP decoders work by solving an approximation of the classical MLD problem known as bit-wise MLD [54]. In the quantum context, an analogous principle is employed: instead of tackling QMLD, we use classical BP decoders and the symplectic representation to solve the problem of qubit-wise MLD. Note that, not only are we not solving QMLD exactly, but we are also ignoring the phenomenon of degeneracy (an optimal quantum decoder would address DQMLD, not QLMD).
Qubit-wise MLD is different from QMLD in that, instead of looking for the most likely error given the syndrome, i.e., looking for the global optimum, we will look for the qubit-wise most likely error, i.e., looking for the marginal optimum. This entails maximizing the probability of each individual qubit given a particular syndrome bit. Qubit-wise MLD can be written as:
@@
where the qubit-wise most likely error is obatined by running through all vaues of $i$: $\hat{E}^{\text{bw}}=[\hat{E}^{\text{bw}}_{1}\ldots\hat{E}^{\text{bw}}_{n}]$. In general, $\hat{E}^{\text{bw}}$ need not coincide with the global optimum obtained by solving (7) [144].
Although computing solutions to (14) is also difficult, this task is amenable to BP. Given a factor graph representation of the parity check matrix of a code, BP can run over the aforementioned graph and obtain the qubit-wise most likely error in polynomial time. Assuming that we are working with CSS codes, note that the rotated planar code belongs to this class, (refer to [144, 54] for a more general description), the BP decoding problem can be understood as the execution of two classical BP decoders, one for $X$-errors and the other for $Z$-errors. Thus, we can describe the iterative schedule of BP decoding as if applied to a classical code [^19] Recall that quantum CSS codes can be understood as the amalgamation of two classical codes, where one code corrects bit-flips and the other corrects phase-flips.. To begin, the BP algorithm requires an error syndrome $s$ and the channel error probability distribution. The error syndrome is computed as $s=H\cdot E$, where $H$ denotes the parity check matrix of the classical code and $E$ represents the error, and is a length nkn-k binary vector. Each of the entries in the syndrome vector is assigned to a check node. Conversely, the channel error probability distribution is made available to the variable nodes. In this sense, the BP decoder operates by executing the following instruction schedule:
@@
    At a given decoding iteration $t$, each variable node, $v_{i}$, sending out a message, $\mu^{t}_{v_{i}\rightarrow c_{j}}$, to the check nodes, $c_{j}$, it is connected to. In the first decoding iteration, $t=1$, this message is equal to the a priori bit error probability, given by the noisy quantum channel in consideration. This message is generally expressed in the log-likelihood domain as
    $\mu^{1}_{v_{i}\rightarrow c_{j}}=l_{\text{ch}}(E_{i})=\text{log}(\frac{p(E_{i}=0)}{p(E_{i}=1)}),$
    where $i\in[1,\ldots,n]$, $j\in[1,\ldots,n-k]$, and the term $l_{\text{ch}}$ represents a priori channel log-likelihood ratio (llr). Note that here we used the slight abuse of notation $E_{i}$, which refers only to a $X$-error or $Z$-error due to the CSS structure of the code, i.e. $E_{i}=E_{i}^{x/z}$. At future iterations, the message sent from a variable node to neighboring check nodes is given by
    $\mu^{t}_{v_{i}\rightarrow c_{j}}=l_{\text{ch}}(E_{i})+\sum_{k=1}^{\sigma-1}\mu^{t-1}_{c_{k}\rightarrow v_{i}},$
    where $\mu^{t-1}_{c_{k}\rightarrow v_{i}}$ are messages received from check nodes in the previous iteration and $\sigma$ is the degree (number of connections) of the variable nodes. Note that the sum goes to $\sigma-1$ because the message received from check node $c_{j}$ in iteration t1t-1 is not considered.
@@
    Once all these initial messages have been received by the check nodes, every check node will reply to all neighbouring variable nodes with the following message:
@@
    where $\psi$ represents the degree of the check nodes, $t$ represents the BP iteration, and $s_{j}$ denotes the syndrome bit associated to that particular check node. Notice how the product required to compute $\mu_{c_{j}\rightarrow v_{i}}$ only goes to $\psi-1$. Once again, this is because the prior variable node message received from the node the current message will be sent to should not be considered.
  3. 3.
    Once all check node messages have been received, variable nodes can compute their llrs (also referred to as beliefs or marginals), which we use to estimate (14). This is done as
    $l^{t}_{\text{ap}}(E_{i})=l_{\text{ch}}(E_{i})+\sum_{k=1}^{\sigma}\mu^{t}_{c_{k}\rightarrow v_{i}}.$
@@
  4. 4.
    At this point we check whether more decoding iterations are needed. First we obtain the estimate $\hat{E}^{\text{bw}}=[\hat{E}^{\text{bw}}_{1}\ldots\hat{E}^{\text{bw}}_{n}]$ by making hard decisions on the beliefs $l^{t}_{\text{ap}}(E_{i})$. These hard decisions commonly follow that negative llr, $l^{t}_{\text{ap}}(E_{i})$, relates to the fact that the estimated probability of error is higher than no error for such qubit. Then, we compute the syndrome associated to the decoding estimate as
    $\hat{s}=H\cdot\hat{E}^{\text{bw}}.$
    If $\hat{s}=s$ then decoding has been successful and the BP algorithm is halted. If not, then additional iterations will be run until either $\hat{s}=s$ or a maximum number of iterations is reached.
@@ 5.3.2 Aftermath of neglecting DQMLD
The excellent performance of BP as a decoding algorithm for classical random-like codes, such as LDPC codes and turbo codes [38, 145], is well-documented. In fact, classical LDPC codes are essentially capacity-achieving when decoded via BP [145]. For this reason, along with their finite-rate guarantees, the design of so-called ‘good’ [^20] By ‘good’ error correction codes we refer to codes whose number of encoded logical qubits and distance increase in a polynomial manner with the number of physical qubits. Essentially, $k,d\approx O(n)$. quantum LDPC codes has been a long-pursued topic in the field of QEC. Although the existence of sparse codes exhibiting such favorable parameter scaling remained unproven for the past two decades, groundbreaking results by Panteleev and Kalachev [48] as well as [146, 147, 148] have finally shown that quantum analogues of robust LDPC codes do actually exist. However, this only addresses half of the problem (the code design aspect), as there are further quantum-specific challenges in terms of decoding. Quantum codes manifest a phenomenon known as degeneracy [54, 128, 67, 149, 144, 80, 46], which has no classical equivalent. This poses a quandary that the classical version of BP cannot resolve, as it is designed for a classical environment in which degeneracy is not present.
@@
Recall that degeneracy in the context of QEC refers to the fact that errors $E\in\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{G}\mkern-1.5mu}\mkern 1.5mu_{n}$ that differ by a stabilizer element (errors that belong to the same logical coset) have the same effect on the code and, thus, are correctable by the same recovery operation. An optimal decoding strategy for degenerate codes should weigh the probability of each logical coset and pick an operator from the most probable one (the operator itself is irrelevant, what matters is picking the right equivalence class). This nuance is critical, as it highlights the differences between optimal decoding for degenerate quantum codes: DQMLD and non-degenerate quantum codes: QMLD. Applying the QMLD rule to a degenerate quantum code is suboptimal, as the operator with highest probability need not belong to the coset with highest probability. This performance difference is further aggravated if a BP decoder is employed to solve QMLD [^21] Recall that BP is an approximation of QMLD that relies on marginalization: it optimizes the error probability individually for each qubit rather than jointly as is called for by QMLD.. This is due to the fact that degenerate quantum codes can simultaneously exhibit an error probability that is sharply peaked over a logical coset (which would lead to great performance under DQMLD) and that has a broad marginal distribution over individual qubits, which given the large number of low and similarly weighted operators of such codes, would severely hinder a BP decoder. The existence of the aforementioned operators and their equivalence under BP decoding lead to the peculiar symmetric degeneracy error phenomenon [144], which is also referred to as split-belief [122] or Quantum Trapping Set (QTS) [150, 151]. Quite frankly, if a code is degenerate enough (if the weight of its generators is substantially smaller than its distance), split beliefs can put the proverbial nail-in-the-coffin for a BP decoder in this context. This cannot be better exemplified than by the fact that that the surface/toric code exhibits no threshold under BP decoding [122]. Surface codes are degenerate by nature, becoming even more so as their size is increased (the weight of their stabilizer generators remains the same but the distance grows). This makes the split-belief phenomenon more prevalent for larger surface codes, explaining the absence of a threshold for this family of codes under BP decoding, as can be seen in FIG. 12.
@@
Split-beliefs have been analyzed in the literature [144, 150, 151] and their impact has been successfully alleviated through myriads of post-processing routines added to general BP decoding. The most successful of these modified BP-based decoding techniques is known as BP-OSD [46]. The performance increments this strategy provides have made it the front-runner in the conversation of a general purpose decoder for QLDPC codes. In fact, the toric code actually exhibits a threshold when decoded via this more sophisticated algorithm. In consequence, this begs the question of whether BP-OSD can compete with the other decoding strategies for the planar code that we have seen thus far.
@@ 5.3.3 Enhanced Belief Propagation: Quantum Ordered Statistics Decoding
The post-processing algorithm known as Ordered Statistics Decoding (OSD) [152, 153] was originally designed to improve the performance of small classical codes, as well as to lower the error floors of certain LDPC codes. It was later adapted to the quantum paradigm by Panteleev and Kalachev in [46], where the authors successfully devised the so called qOSD routine via specific modifications to the classical OSD algorithm. In a similar fashion to other post-processing routines for sparse quantum codes, qOSD only works in conjunction with a BP decoder; i.e., it requires the soft outputs of a BP decoder [^22] Or any other decoder capable of yielding soft values as its output. Soft in this context means that the hard decisions on the most probable error have not been taken so essentially it is aprobability distribution. in order to function. For this reason, the decoding routine that combines both BP and qOSD is generally referred to as BP-OSD. In later work, BP-OSD was shown to work well for Toric codes and a novel class of semi-topological codes [122], and recently, it has also been shown to be a valid decoding strategy for bias-tailored QLDPC codes [154]. The authors of [122] also made their implementation of BP-OSD public, which marked the first open-source demonstration of this decoding algorithm.
@@
As is done in [122], for the sake of notational simplicity, we describe OSD post-processing as applied to the classical decoding problem
$s=H\cdot e.$ (15)
@@
It is obvious that this framework is equally applicable (albeit with minor programming modifications) to decoding the $X$ and $Z$ components of a CSS code or directly decoding over the entire parity check matrix of a non-CSS code.
@@
The OSD post-processing algorithm is called upon whenever BP fails to produce an estimate $\hat{e}$ that matches the measured syndrome. Hence, our starting point is the following: after attempting to decode the measured syndrome $s$ via BP we end up with the incorrect estimate of the error $\hat{e}$. The estimate is known to be erroneous since its syndrome does not match the true syndrome. However, the fact that $s\neq H\cdot\hat{e}$ does not imply that all components of $\hat{e}$ are wrong, a rationale that OSD will exploit to find a valid solution to (15).
@@
We begin by introducing a necessary set of concepts. We will call the set of indices $[I]$ for which $H\cdot e_{I}=H\cdot\hat{e}_{I}$ the most reliable information set. The complement of this set $\bar{I}]$, all those indices for which $H\cdot e_{\bar{I}}\neq H\cdot\hat{e}_{\bar{I}}$, will thus be referred to as the least reliable information set. As we explain in what follows, OSD post-processing exploits the concept of reliable information sets to solve the linear system described by (15). The parity check matrix $H$ of a quantum error correction code is a rectangular matrix that does not have full column-rank, making it impossible to solve $s=H\cdot\hat{e}$ via matrix inversion. However, the system described by an appropriate set of nkn-k linearly independent columns $[S]$ of $H$ is actually solvable:
$s=H_{[S]}e_{[S]},$ (16)
@@
$H^{-1}_{[S]}s=e_{[S]},$ (17)
(note that $H_{[S]}$ is a full-rank matrix). For every choice of $[S]$, the basis of linearly independent columns, we will obtain a unique solution $e_{[S]}$, that satisfies (16). Against this backdrop, we can understand what OSD post-processing is about: if a full-column rank subset of the parity check matrix can be found and used to solve (16) via matrix inversion, we will always find a solution $e_{[S]}$ that produces a matching syndrome $s$. Additionally, because this solution is unique, any symmetries that might hinder BP decoding (like those that cause split beliefs) are now broken.
@@
OSD-0: Assume the following, after measuring a given syndrome $s$ for a specific code with parity check matrix $H$, we have decoded via BP and obtained an estimate of the error $\hat{e}$, which unfortunately does not produce a matching syndrome: $s\neq H\cdot\hat{e}$. In this context, we would execute OSD, which would operate as follows:
@@
    Take the soft-outputs [^24] The a-posteriori llrs estimated in the final decoding iteration. of the BP decoder given by $l_{i,\text{ap}}=\frac{P(e_{i}=0|\mathbf{s})}{P(e_{i}=1|\mathbf{s})},\ \forall i\in{0,\ldots,n}$, and order them from most-likely to least-likely to have been flipped (increasing order of magnitude). Store the list of bit-indices $\text{OG}]$, as this defines the least reliable information set of bits.
@@
    Invert $\Lambda_{[J]}$ into $\Lambda^{-1}_{[J]}$. Calculate the solution $e_{[J]}$ to the OSD syndrome equation as $\Lambda^{-1}_{[J]}s=e_{[J]}$.
@@
    The complete solution to the decoding problem is $\mathbf{e}_{\Lambda_{0}}=[e_{[J]},e_{[\bar{J}]}]$, where $J$ and $\bar{J}$ denote the least and most reliable information set of bits, respectively. Knowing that $e_{[J]}$ satisfies $\Lambda_{[J]}e_{[J]}=s$ then it is easy to see that $\Lambda\mathbf{e}_{\Lambda_{0}}=\Lambda[e_{[J]},\mathbf{0}]=s$.
@@
    The last step is to take the solution $\mathbf{e}_{\Lambda_{0}}=[e_{[J]},\mathbf{0}]$, and map it to the original bit-index order: $\mathbf{e}_{\Lambda_{0}}\rightarrow\mathbf{e}_{\text{OSD-0}}$. We call the rearranged vector, $e_{\text{OSD-0}}$, the OSD-0 solution.
@@
In FIG. 13, an example is provided where a specific error produces a syndrome which is later decoded finally obtaining a recovered error. In the top row, the error and the syndrome within the surface code are introduced, data qubits, $X$ and $Z$-checks are labelled in black, green and orange colors respectively. In the second row, one can see a graphical representation of the BP method, where messages are sent from data qubits to checks and vice versa. Notice, how there are two grey circles for each data qubit label, that is because the two BP graphs are independent, and so each data qubit will have two resulting marginal probabilities, one for the $X$-operators and another one for the $Z$ ones. In the following row, the resulting marginal probabilities are represented within the data qubits of the surface code with $X$-check qubits and $Z$-check ones. On the left surface code a data qubit being redder indicate its marginal probability of recovering an $X$-operator being larger, and thus its llr being lower, and whiter otherwise. The same rules go by the $Z$-check surface code on the right. On the following row, the matrix $\Lambda$ is represented. The chosen data qubits are the ones which represent the independent columns with the higher marginal probabilities, the four columns on the left are extracted from the $X$-check subgraph and the remaining four from the $Z$-check subgraph, as indicated by the labels on top of the matrix. This matrix is inverted following the earlier described process, reaching a recovered error which is depicted in the last row. Notice how the error is not the same set of Pauli operators as the inputted one, nevertheless, it is within the same stabilising group, and thus, successfully corrects the error.
Figure 13: Graphical description of a BPOSD-0 process for a specific syndrome in a 3x3 surface code.

### v5

@@ 5.3.1 BP: Specifics
In the classical context, BP decoders work by solving an approximation of the classical maximum likelihood decoding problem known as bit-wise maximum likelihood decoding [41]. In the quantum context, an analogous principle is employed: instead of tackling QMLD, we use classical BP decoders to solve the problem of qubit-wise maximum likelihood decoding.
Qubit-wise maximum likelihood decoding is different from QMLD in that, instead of looking for the most likely error given the syndrome, we will look for the qubit-wise most likely error, i.e., looking for the marginal probabilities of every column within the parity check matrix. Qubit-wise maximum likelihood decoding can be written as:
@@
where the qubit-wise most likely error is obtained by running through all values of $i$: $\hat{E}^{\text{bw}}=[\hat{E}^{\text{bw}}_{1}\ldots\hat{E}^{\text{bw}}_{n}]$. In general, $\hat{E}^{\text{bw}}$ need not coincide with the global optimum obtained by solving (7) [118].
Although computing solutions to (23) is also difficult, this task is amenable to BP. Given the Tanner graph of a code, BP can run over the aforementioned graph and obtain the qubit-wise most likely error in polynomial time. As earlier seen, we can consider the code as two separate parity check matrices $H_{X}$ and $H_{Z}$ and, thus, the BP decoding problem can be understood as the execution of two classical BP decoders that are run simultaneously, one for $X$-errors and the other for $Z$-errors. Thus, to understand BP-based decoding of CSS codes, we can limit ourselves to explaining how BP decoding works for classical codes To begin, the BP algorithm requires the following:
  • A Tanner graph $G$ representing the parity check matrix $H$ of the $[n,k,d]$ classical code [^25] The $[n,k,d]$ notation represents the fact that the code has $n$ physical bits, $k$ logical bits, and distance $d$. we are working with.
  • An error syndrome $\bar{s}$ computed as $\bar{s}=H\cdot\bar{e}$.
  • The probability distribution of the error, P(e^)P($\hat${e}).
Once $G$, $\bar{s}$, and P(e^)P($\hat${e}) have been computed, the BP decoding algorithm operates in the following manner:
@@
    Initialization: Each entry in the syndrome vector $\{s\}_{1}^{m}$ is provided to a check node $\{c\}_{1}^{m}$ of $G$, where $m$ indicates the number of checks. Additionally, each variable node $\{v\}_{1}^{n}$ is assigned a probability, known as the a priori bit error probability, computed as
    $\frac{p(\hat{e}_{i}=0)}{p(\hat{e}_{i}=1)},$
    where $i\in[1,\ldots,n]$. Once this information is provided to the check and variable nodes of $G$ the BP algorithm begins exchanging messages between them. These messages are understood in the context of decoding iterations, which are defined as the bi-directional exchange of messages over all the edges of $G$. This means that an iteration completes whenever all connected variable and check nodes have both sent and received a message. The BP algorithm operates by executing numerous of these iterations to propagate information throughout $G$. In what follows we describe the message exchange procedure for an arbitrary iteration.
@@
    At any given decoding iteration $t$, each variable node, $v_{i}$, sends out a message, $\mu^{t}_{v_{i}\rightarrow c_{j}}$, to the check nodes, $c_{j}$, it is connected to. In the very first decoding iteration, $t=1$, this message is equal to the a priori bit error probability which was assigned to each variable node in the initialization stage. These messages are generally expressed in the log-likelihood domain. For $t=1$ the messages are given by
    $\mu^{1}_{v_{i}\rightarrow c_{j}}=l_{\text{ch}}(\hat{e}_{i})=\text{log}(\frac{p(\hat{e}_{i}=0)}{p(\hat{e}_{i}=1)}),$
    where $i\in[1,\ldots,n]$, $j\in[1,\ldots,n-k]$, and the term $l_{\text{ch}}$ represents the a priori channel log-likelihood ratio (llr). In future iterations ($t>1$), the message sent from a variable node to neighboring check nodes is given by
    $\mu^{t}_{v_{i}\rightarrow c_{j}}=l_{\text{ch}}(\hat{e}_{i})+\sum_{k=1}^{\sigma-1}\mu^{t-1}_{c_{k}\rightarrow v_{i}},$
    where $\mu^{t-1}_{c_{k}\rightarrow v_{i}}$ are messages received from check nodes in the previous iteration and $\sigma$ is the degree (number of connections) of the variable nodes. The sum goes to $\sigma-1$ because the message received from check node $c_{j}$ in the previous iteration (t1t-1) is not considered.
  3. 3.
    Once the check nodes have received a message from neighboring variable nodes, every check node will reply to all neighbouring variable nodes with the following message:
@@
    where $\psi$ represents the degree of the check nodes, and $s_{j}$ denotes the syndrome bit associated to that particular check node. Notice how the product required to compute $\mu_{c_{j}\rightarrow v_{i}}$ only goes to $\psi-1$. Once again, this is because the prior variable node message received from the node the current message will be sent to should not be considered.
  4. 4.
    Once all check node messages have been received by the variable nodes, the message exchange process for iteration $t$ has finished. At this point, variable nodes update their original probabilities (also referred to as beliefs or marginals) based on the information transmitted from the check nodes. Given this update, the a-priori llrs now become a-posteriori llrs. This update is done as
    $l^{t}_{\text{ap}}(\hat{e}_{i})=l_{\text{ch}}(\hat{e}_{i})+\sum_{k=1}^{\sigma}\mu^{t}_{c_{k}\rightarrow v_{i}}.$
@@
  5. 5.
    These updated a posteriori llrs is what the BP decoder uses to estimate (23). First we obtain the estimate $\hat{e}$ by making hard decisions on the beliefs, $l^{t}_{\text{ap}}(\hat{e}_{i})$ , that is, we choose a threshold value for $l^{t}_{\text{ap}}(\hat{e}_{i})$ at which we consider the parity to be non-trivial. The common threshold consists in considering $l^{t}_{\text{ap}}(\hat{e}_{i})<0$ to be flipped parities. Then, we compute the syndrome associated to the decoding estimate as
    $\hat{s}=H\cdot\hat{e}$.
    If $\hat{s}=\bar{s}$ then decoding has been successful and the BP algorithm is halted. If not, then additional iterations will be run until either $\hat{s}=\bar{s}$ or a maximum number of iterations is reached.
Therefore, one can see the BP algorithm as a method in which both variable and check nodes iterative send information about their state to their adjacent nodes regarding their marginal probability state. After sending said information in the form of a belief, their state is updated through the beliefs they receive. This local transmission of information which encompasses all the Tanner graph produces changes in the marginal probabilities of the variable nodes until it is halted. Afterwards, one can retrieve a recovered error $\hat{e}$ by considering that negative $l^{t}_{\text{ap}}(\hat{e}_{i})$ correspond to flipped parities ($\hat{e}_{i}=1$). This recovery is taken under the consideration that negative $l^{t}_{\text{ap}}(\hat{e}_{i})$ mean resulting marginal probabilities indicating the variable node is more likely to be flipped than not to.
@@ 5.3.2 Aftermath of neglecting DQMLD
The excellent performance of BP as a decoding algorithm for classical random-like codes, such as LDPC codes and turbo codes [119], is well-documented. In fact, classical LDPC codes essentially achieve theoretical performance limits when decoded via BP [119]. For this reason, along with their finite-rate guarantees, the design of so-called ‘good’ [^26] By ‘good’ error correction codes we refer to codes whose number of encoded logical qubits and distance increase in a polynomial manner with the number of physical qubits. Essentially, $k,d\approx\mathcal{O}(n)$. quantum LDPC codes has been a long-pursued topic in the field of QEC. Although the existence of sparse codes exhibiting such favorable parameter scaling remained unproven for the past two decades, groundbreaking results by Panteleev and Kalachev [38] as well as [120, 121, 122] have finally shown that quantum analogues of robust LDPC codes do actually exist. In light of these results and given the overhead advantages that quantum LDPC codes offer when compared to surface codes, industry leading companies have begun studying their commercial viability, as was shown in recently published results for circuit-level performance of quantum LDPC-based memories [123, 124].
@@
The design of good quantum LDPC codes is only half the problem, however, as decoding quantum codes with the classical BP algorithm leads to performance that is far from what would be expected based on the classical literature. Previously in this work we highlighted the differences between optimal decoding for degenerate quantum codes: DQMLD and non-degenerate quantum codes: QMLD. Applying the QMLD rule to a degenerate quantum code is suboptimal, as the operator with highest probability need not belong to the coset with highest probability. Quantum codes, especially quantum LDPC codes, are highly degenerate, so a performance penalty is already incurred when settling for the QMLD rule. This performance difference is further aggravated if a BP decoder is chosen to solve QMLD [^27] Recall that BP is an approximation of QMLD that relies on marginalization: it optimizes the error probability individually for each qubit rather than jointly as is called for by QMLD.. This is due to the fact that degenerate quantum codes can simultaneously exhibit an error probability that is sharply peaked over a logical coset (which would lead to great performance under DQMLD) and that has a broad marginal distribution over individual qubits, which given the large number of low and similarly weighted operators of such codes, would severely hinder a BP decoder. The existence of the aforementioned operators and their equivalence under BP decoding lead to the peculiar symmetric degeneracy error phenomenon [118], which is also referred to as split-belief [98] or Quantum Trapping Set (QTS) [125]. Quite frankly, if a code is degenerate enough (if the weight of its generators is substantially smaller than its distance), split beliefs can put the proverbial nail-in-the-coffin for a BP decoder in this context. This cannot be better exemplified than by the fact that the rotated code exhibits no threshold under BP decoding [98]. Rotated codes are degenerate by nature, becoming even more so as their size is increased (the weight of their stabilizer generators remains the same but the distance grows). This makes the split-belief phenomenon more prevalent for larger planar codes, explaining the absence of a threshold for this family of codes under BP decoding, as can be seen in FIG. 13.
@@
Split-beliefs have been analyzed in the literature [118, 125] and their impact has been successfully alleviated through myriads of post-processing routines added to general BP decoding. The most successful of these modified BP-based decoding techniques is known as BP-OSD [37]. The performance increments this strategy provides have made it the front-runner in the conversation of a general purpose decoder for QLDPC codes. In fact, the rotated code actually exhibits a threshold when decoded via this more sophisticated algorithm. In consequence, this begs the question of whether BP-OSD can compete with the other decoding strategies for the planar code that we have seen thus far.
@@ 5.3.3 Enhanced Belief Propagation: Quantum Ordered Statistics Decoding
The post-processing algorithm known as Ordered Statistics Decoding (OSD) [126, 127] was originally designed to improve the performance of small classical codes, as well as to lower the error floors of certain LDPC codes. It was later adapted to the quantum paradigm by Panteleev and Kalachev in [37], where the authors successfully devised the so called qOSD routine via specific modifications to the classical OSD algorithm. In a similar fashion to other post-processing routines for sparse quantum codes, qOSD only works in conjunction with a BP decoder; i.e., it requires the marginal probability outputs of a BP decoder [^28] Or any other decoder capable of yielding soft values as its output. Soft in this context means that the hard decisions on the most probable error have not been taken so essentially it is a probability distribution. in order to function. For this reason, the decoding routine that combines both BP and qOSD is generally referred to as BP-OSD. In later work, BP-OSD was shown to work well for Toric codes and a novel class of semi-topological codes [98], and recently, it has also been shown to be a valid decoding strategy for bias-tailored QLDPC codes [128].
@@
As is done in [98], for the sake of notational simplicity, we describe OSD post-processing as applied to the classical decoding problem described by
$\bar{s}=H\cdot\bar{e}.$ (24)
@@
As has previously seen, it is obvious that this framework is equally applicable (albeit with minor programming modifications) to decoding the $X$ and $Z$ components of a CSS code.
@@
The OSD post-processing algorithm is called upon whenever BP fails to produce an estimate $\hat{e}$ that matches the measured syndrome. Hence, our starting point is the following: after attempting to decode the measured syndrome $\bar{s}$ via BP we end up with the incorrect estimate of the error $\bar{s}$. The estimate is known to be erroneous since its syndrome does not match the true syndrome. However, the fact that $\bar{s}\neq H\cdot\hat{e}$, where $H$ is the parity check matrix does not imply that all components of $\hat{e}$ are wrong, a rationale that OSD will exploit to find a valid solution to (24).
@@
We begin by introducing a necessary set of concepts. We will call the set of indices $[I]$ for which $H\cdot\bar{e}_{I}=H\cdot\hat{e}_{I}$ the most reliable information set. The complement of this set $\bar{I}]$, all those indices for which $H\cdot\bar{e}_{\bar{I}}\neq H\cdot\hat{e}_{\bar{I}}$, will thus be referred to as the least reliable information set. As we explain in what follows, OSD post-processing exploits the concept of reliable information sets to solve the linear system described by (24). The parity check matrix $H$ of a quantum error correction code is a rectangular matrix that does not have full column-rank, making it impossible to solve $\bar{s}=H\cdot\bar{e}$ via matrix inversion. However, the system described by an appropriate set of nkn-k linearly independent columns $[S]$ of $H$ is actually solvable:
$\bar{s}=H_{[S]}\hat{e}_{[S]},$ (25)
@@
$H^{-1}_{[S]}\bar{s}=\hat{e}_{[S]},$ (26)
(note that $H_{[S]}$ is a full-rank matrix). For every choice of $[S]$, the basis of linearly independent columns, we will obtain a unique solution $\hat{e}_{[S]}$, that satisfies (25). Against this backdrop, we can understand what OSD post-processing is about: if a full-column rank subset of the parity check matrix can be found and used to solve (25) via matrix inversion, we will always find a solution $\hat{e}_{[S]}$ that produces a matching syndrome $\bar{s}$. Additionally, because this solution is unique, any symmetries that might hinder BP decoding (like those that cause split beliefs) are now broken.
@@
OSD-0: Assume the following, after measuring a given syndrome $\bar{s}$ for a specific code with parity check matrix $H$, we have decoded via BP and obtained an estimate of the error $\hat{e}$, which unfortunately does not produce a matching syndrome: $\bar{s}\neq H\cdot\hat{e}$. In this context, we would execute OSD, which would operate as follows:
@@
    Take the soft-outputs [^30] The a-posteriori llrs are estimated in the final decoding iteration. of the BP decoder given by $l_{i,\text{ap}}=\frac{P(\hat{e}_{i}=0|\bar{s})}{P(\hat{e}_{i}=1|\bar{s})},\ \forall i\in{0,\ldots,n}$, and order them from most-likely to least-likely to have been flipped (increasing order of magnitude). Store the list of bit-indices $\text{OG}]$, as this defines the least reliable information set of bits.
@@
    Invert $\Lambda_{[J]}$ into $\Lambda^{-1}_{[J]}$. Calculate the solution $\hat{e}_{[J]}$ to the OSD syndrome equation as $\Lambda^{-1}_{[J]}\bar{s}=\hat{e}_{[J]}$.
@@
    The complete solution to the decoding problem is $\hat{e}_{\Lambda_{0}}=[\hat{e}_{[J]},\hat{e}_{[\bar{J}]}]$, where $J$ and $\bar{J}$ denote the least and most reliable information set of bits, respectively. Knowing that $\hat{e}_{[J]}$ satisfies $\Lambda_{[J]}\hat{e}_{[J]}=\bar{s}$ then it is easy to see that $\Lambda\hat{e}_{\Lambda_{0}}=\Lambda[\hat{e}_{[J]},\mathbf{0}]=\bar{s}$.
@@
    The last step is to take the solution $\mathbf{e}_{\Lambda_{0}}=[e_{[J]},\mathbf{0}]$, and map it to the original bit-index order: $\hat{e}_{\Lambda_{0}}\rightarrow\hat{e}_{\text{OSD-0}}$. We call the rearranged vector, $e_{\text{OSD-0}}$, the OSD-0 solution.
@@
In FIG. 14, an example is provided where a specific error produces a syndrome which is later decoded finally obtaining a recovered error. In FIG. 14 a), the error and the syndrome within the rotated code are introduced, data qubits, $X$ and $Z$-checks are labelled in black, green and orange colors respectively. In FIG. 14 b), one can see a graphical representation of the BP method, where messages are sent from data qubits to checks and vice versa. Notice, how there are two white circles for each data qubit label, that is because the two BP graphs are independent, and so each data qubit will have two resulting marginal probabilities, one for the $X$ operators and another one for the $Z$ ones. In FIG. 14 c) and d), the resulting marginal probabilities are represented within the data qubits of the rotated code with $X$-check qubits and $Z$-check ones. Data qubit being redder indicate its marginal probability of recovering a non-trivial operator being larger, and thus its llr being lower, and whiter otherwise. In FIG. 14 e), the matrix $\Lambda$ is represented. The chosen data qubits are the ones which represent the independent columns with the higher marginal probabilities, the four columns on the left are extracted from the $X$-check subgraph and the remaining four from the $Z$-check subgraph, as indicated by the labels on top of the matrix. This matrix is inverted following the earlier described process, reaching a recovered error which is depicted In FIG. 14 f). Notice how the error is not the same set of Pauli operators as the inputted one, nevertheless, it is within the same stabilising group, and thus, successfully corrects the error.
Figure 14: Graphical description of a BPOSD-0 process for a specific syndrome in a distance-3 rotated code in a). The process begins by running BP over the Tanner graph in b) resulting in the marginal probabilities in c) and d) (lower llr implies redder color). The parity check matrix is ordered according to the llrs. The invertible matrix associated to the least reliable set is selected in e). Gaussian elimination is done with the obtained $\Lambda$ matrix to solve the syndrome problem, leading to a recovered error matching the syndrome in f).